OpenStudy (anonymous):

Find any critical numbers of the function. (4x)/(x^2+1)

OpenStudy (anonymous):

\[\frac{ 4x }{ x ^{2}+1}\] is the equation.

OpenStudy (anonymous):

what you want?

OpenStudy (anonymous):

We are working on applications of differentiation, specifically extremums. What I learned is you first take the derivative, then figure out when it equals zero or is undefined. The numbers where it equal zero or is undefined is what I'm looking for.

OpenStudy (anonymous):

you want to i make drive of your equation?

OpenStudy (anonymous):

(u/v)' = (u'v-v'u)/(v^2)

OpenStudy (anonymous):

by get u=4x and v=x^2+1 solve it

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Well actually I did that, but you just made me think of what I forgot to do and why it's all messed up- I forgot the 'du' of the derivative part of 'x^2 +1'! Thanks! Haha

OpenStudy (anonymous):

Wait, I lied.... it doesn't need a "du". I'm not sure what I did wrong still to get the bizarre results that I got... Can you go through all of it in steps for me?

OpenStudy (anonymous):

I know how to do the derivative, I just think that maybe somewhere along the lines I made a mistake that I'm having trouble catching because the ending equations that I got were just incredibly weird. The only way I could possibly solve them is with i.

OpenStudy (anonymous):

( 4(x^2+1)-2x(4x) )/( (x^2+1)^2)

OpenStudy (anonymous):

Okay, I've got that part.

OpenStudy (anonymous):

Next, I have simplified it to \[\frac{ -4x ^{2}+4 }{ (x^{2}+1)^{2} }\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, that's good, but when I set the bottom to equal zero- like (x^2+1)^2=0 - it ends up to x^2=-1, and you can't take a square root of -1. :/ So I'm not sure what's up with that.

OpenStudy (anonymous):

x^2=-1 x=i=sqr(-1)

OpenStudy (anonymous):

But I don't think that we can use i, so that's what I'm wondering about.

OpenStudy (anonymous):

if you write question correctly the answer is x=i

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

your welcome